. A baseball player hits a ball such that as it leaves the bat, it has speed of
40 meters per second and has an angle of 44 degrees
. What is the time that the ball
reaches the highest point of its flight?
1
Expert's answer
2021-08-29T16:57:00-0400
"v_y=v_{0y}-gt,""0=v_0sin\theta-gt,""t=\dfrac{v_0sin\theta}{g}=\dfrac{40\ \dfrac{m}{s}\cdot sin44^{\circ}}{9.8\ \dfrac{m}{s^2}}=2.83\ s."